### v1

# Perspectives on the arithmetic nature of the ratios $\zeta(2n+1)/\pi^{2n+1}$ and $\beta(2n)/\pi^{2n}$

Abstract: We investigate the values of the Riemann zeta function at odd integers and the Dirichlet beta function at even integers, by collecting several distinct analytic frameworks converging to these values, thus providing a unifying perspective. Beyond analytic interest, these formulas motivate linear independence conjectures which, if established, would imply the irrationality of the quantities $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$.
Introduction
Euler is renowned for his resolution of the Basel problem, establishing that
$\zeta(2)=\frac{\pi^{2}}{6}.$
Moreover, he derived a general formula for the values of the Riemann zeta function at even positive integers, namely
$\zeta(2n)=(-1)^{\,n+1}\,\frac{B_{2n}\,(2\pi)^{2n}}{2\,(2n)!},$
where $B_{2n}$ denotes the $2n$-th Bernoulli number [1]. Despite his extensive efforts, Euler was unable to obtain an analogous formula for $\zeta(2n+1)$. Since then, the arithmetic nature of these odd zeta values has remained a central open problem, attracting sustained attention from mathematicians to this day.
A closely related, though less widely studied, function is Dirichlet’s beta function, defined for $\Re(s)>0$ by
$\beta(s):=\sum_{m=0}^{\infty}\frac{(-1)^{m}}{(2m+1)^{s}}.$
Although this $L$-function shares many structural similarities with $\zeta(s)$, it has received comparatively limited attention in the literature. Historically, even prior to Euler’s investigations, Leibniz demonstrated the identity
$\sum_{n=0}^{\infty}\frac{(-1)^{n}}{2n+1}\;=\;\frac{\pi}{4}.$
In contrast to the Riemann zeta function, whose explicit values are known at even arguments but remain mysterious at odd ones, the Dirichlet beta function exhibits a complementary phenomenon. For odd arguments, one has explicit closed‐form evaluations [2]:
$\beta(2n+1)=(-1)^{n}\,\frac{E_{2n}\,\pi^{2n+1}}{2^{2n+2}\,(2n)!},$
where $E_{2n}$ are Euler numbers. Hence, the ratios $\beta(2n+1)/\pi^{2n+1}$ are rational, underscoring a striking parallel with the special values of $\zeta(s)$. For even arguments of $\beta(s)$, however, no such closed formula is currently known.
According to Euler’s formula for $\zeta(2n)$ and the expression of $\beta(2n+1)$, many tried to find some rational coefficients $r_{n}$ and $s_{n}$ such that the relations may be satisfied
$\zeta(2n+1)=r_{n}\,\pi^{2n+1}\quad\text{and }\quad\beta(2n)=s_{n}\,\pi^{2n}$
Given the futility of the extraordinarily numerous attempts to compute these numbers, the solution of the designated problem is generally regarded as impossible; but so far a strong proof of this impossibility is lacking. Only heuristic methods and deep computations by brute force support this impossibility to this day. One may solve the problem by proving that the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ are not multiples of $\pi$. The difficulty of this new reformulation partly stems from the lack of sufficiently many closed-form expressions or identities involving these values that might serve as a starting point for an irrationality or transcendence proof. This paper is devoted to some integral representations of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ through interesting logarithmic, trigonometric and hyperbolic kernels, along with several conjectures that support their non representability as rational multiples of $\pi$.
§1. Preliminaries
The point of departure of our investigations was the Malmsten’s integral
$I=\displaystyle\int_{0}^{1}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx$
that has been firstly evaluated to $\dfrac{7\zeta(3)}{8\pi^{2}}$ by Iaroslav V, Blagouchine in 2014 [4]. Therefore the proof of its convergence is omitted here for the sake of conciseness. Notwithstanding, the presence of singularities at its bounds requires to treat it with some care. As consequence, we rewrite the integral as the limit $I=\displaystyle\lim_{\xi\to 0^{+}}\int_{\xi}^{1-\xi}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx$
We consider first $u^{\prime}(x)=\dfrac{x(x^{4}-4x^{2}+1)}{(1+x^{2})^{4}}$ and we show easily that its anti-derivative is $u(x)=\dfrac{x^{2}(1-x^{2})}{2(1+x^{2})^{3}}$. Then we consider $v(x)=\ln\ln\frac{1}{x}$ and observe that its derivative is $v^{\prime}(x)=\dfrac{1}{x\ln x}$. Since both functions $u,v\in\mathcal{C}^{1}[\xi,1-\xi]$, an integration by parts yields
$I=\displaystyle\lim_{\xi\to 0^{+}}\left(\left[\dfrac{x^{2}(1-x^{2})\ln\ln\frac{1}{x}}{2(1+x^{2})^{3}}\right]_{\xi}^{1-\xi}-\dfrac{1}{2}\int_{\xi}^{1-\xi}\dfrac{x^{2}(1-x^{2})}{(1+x^{2})^{3}x\ln x}dx\right)$
Which implies
$I=\displaystyle\underbrace{\lim_{\xi\to 0^{+}}\left[\dfrac{x^{2}(1-x^{2})\ln\ln\frac{1}{x}}{2(1+x^{2})^{3}}\right]_{\xi}^{1-\xi}}_{0}-\lim_{\xi\to 0^{+}}\dfrac{1}{2}\int_{\xi}^{1-\xi}\dfrac{x^{2}(1-x^{2})}{(1+x^{2})^{3}x\ln x}dx$
It remains
$I=\displaystyle\lim_{\xi\to 0^{+}}\dfrac{1}{2}\int_{\xi}^{1-\xi}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}dx=\int_{0}^{1}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}\,dx$
The change of variable $x\rightarrow\tan u$ is proceeded in this manner :
$\displaystyle\int_{0}^{1}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}\,dx=\dfrac{7\zeta(3)}{4\pi^{2}}$ $\Rightarrow\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\tan u(\tan^{2}u-1)}{(1+\tan^{2}u)^{3}\ln(\tan u)}(1+\tan^{2}u)\,du=\dfrac{7\zeta(3)}{4\pi^{2}}$
$\Rightarrow\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\cos^{4}u\tan u(\tan^{2}u-1)}{\ln(\tan u)}\,du=\dfrac{7\zeta(3)}{4\pi^{2}}$
$\Rightarrow\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\cos u\sin u(\sin^{2}u-\cos^{2}u)}{\ln(\tan u)}\,du=\dfrac{7\zeta(3)}{4\pi^{2}}$
$\Rightarrow-\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\cos u\sin u\cos(2u)}{\ln(\tan u)}\,du=\dfrac{7\zeta(3)}{4\pi^{2}}$
$\Rightarrow\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4u}{\ln(\tan u)}\,du=-\dfrac{7\zeta(3)}{\pi^{2}}\quad(1.1)$
Blagouchine introduced in fact more generally the family of integrals on page 80:
$I_{n}=\int_{0}^{1}\frac{x^{n-1}\ln\ln\frac{1}{x}}{(1+x^{2})^{n}}\,dx=\int_{1}^{\infty}\frac{x^{n-1}\ln\ln x}{(1+x^{2})^{n}}\,dx=\frac{1}{2n}\int_{0}^{\infty}\frac{\ln x}{\cosh^{n}x}\,dx,$
And with contour integration technique, he obtained further[4] :
$I_{2}=-\frac{1}{2}\ln 2+\frac{1}{4}\ln\pi-\frac{\gamma}{4}$
$I_{6}=-\frac{1}{60}\ln 2+\frac{1}{120}\ln\pi-\frac{\gamma}{120}-\frac{7\,\zeta(3)}{192\pi^{2}}-\frac{31\,\zeta(5)}{320\pi^{4}}$
where $\gamma$ denotes the Euler-Mascheroni constant. We can rewrite $I_{6}$ as
$I_{6}=-\frac{1}{30}I_{2}-\frac{7\,\zeta(3)}{192\pi^{2}}-\frac{31\,\zeta(5)}{320\pi^{4}}$
And by replacing $I_{6},I_{2}$ and $\zeta(3)$ by their $\ln\ln$ integral representations, it follows:
$\int_{0}^{1}\dfrac{x(x^{8}-26x^{6}+66x^{4}-26x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{6}}\,dx=-\dfrac{93\zeta(5)}{8\pi^{4}}$
Same remarks and same algebra on $I_{8}$ yield:
$\int_{0}^{1}\dfrac{x(x^{12}-120x^{10}+1\,191x^{8}-2\,416x^{6}+1\,191x^{4}-120x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{8}}\,dx=\dfrac{5\,715\zeta(7)}{16\pi^{6}}$
And by pursuing the same line of reasoning as above, the two later Malmsten’s integrals take the forms :
$\dfrac{\zeta(5)}{\pi^{4}}=-\dfrac{1}{186}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4x}{\ln(\tan x)}\,dx+\dfrac{1}{124}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 8x}{\ln(\tan x)}\,dx\quad(1.2)$
$\frac{\zeta(7)}{\pi^{6}}=-\frac{17}{91\,440}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4x)}{\ln(\tan x)}\,dx+\frac{1}{1\,524}\int_{0}^{\frac{\pi}{4}}\frac{\sin(8x)}{\ln(\tan x)}\,dx-\frac{1}{2\,032}\int_{0}^{\frac{\pi}{4}}\frac{\sin(12x)}{\ln(\tan x)}\,dx\quad(1.3)$
These results are highlighting general patterns. In fact, it is easy to recognize the pattern
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(4nx)}{\ln(\tan x)}\,dx$
where $n\in\mathbb{N}$.
§2. Main Result
Deeper linear algebra through Gauss elimination algorithm on (1.1), (1.2) and (1.3) allows to conclude following :
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(4x)}{\ln(\tan x)}\,dx=-7\dfrac{\zeta(3)}{\pi^{2}}$
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(8x)}{\ln(\tan x)}\,dx=-\dfrac{14}{3}\dfrac{\zeta(3)}{\pi^{2}}+124\dfrac{\zeta(5)}{\pi^{4}}$
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(12x)}{\ln(\tan x)}\,dx=-\dfrac{161}{45}\dfrac{\zeta(3)}{\pi^{2}}+\dfrac{496}{3}\dfrac{\zeta(5)}{\pi^{4}}-2\,032\dfrac{\zeta(7)}{\pi^{6}}$
After some resilience and tact, the following equalities also hold true :
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 16x}{\ln(\tan x)}\,dx=-\dfrac{44}{15}\dfrac{\zeta(3)}{\pi^{2}}+\dfrac{2\,728}{15}\dfrac{\zeta(5)}{\pi^{4}}-4\,064\dfrac{\zeta(7)}{\pi^{6}}+32\,704\dfrac{\zeta(9)}{\pi^{8}}$
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 20x}{\ln(\tan x)}\,dx=-\dfrac{563}{225}\dfrac{\zeta(3)}{\pi^{2}}+\dfrac{178\,064}{945}\dfrac{\zeta(5)}{\pi^{4}}-\dfrac{87\,376}{15}\dfrac{\zeta(7)}{\pi^{6}}+\dfrac{261\,632}{3}\dfrac{\zeta(9)}{\pi^{8}}-524\,032\dfrac{\zeta(11)}{\pi^{10}}$
On this basis, we are led to state the following conjecture :
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(4nx)}{\ln(\tan x)}\,dx=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}$
The proof of this statement is fully detailed in [3] and is omitted here in the interest of succinctness. We merely note the existence of integrals $I_{n}$ satisfying
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}\quad(2.4)$
Furthermore, we also find in [3] some integrals $I_{n}$ which adhere to this form
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\beta(2p)}{\pi^{2p-1}}\quad(2.5)$
In fact, the existence of integrals holding either the form (2.4) or (2.5) is the main and central result of this paper. The following table summarizes some of them, where $n\in\mathbb{N}^{*}$ and $C_{p,n}$ being rational coefficients with $p$ running over the set $\llbracket 1,n\rrbracket$
$\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}$ $\sum_{p=1}^{n}C_{p,n}\,\frac{\beta(2p)}{\pi^{2p-1}}$
$\tiny{\displaystyle\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx\quad(k\in\llbracket 0,n-1\rrbracket)}$ $\int_{0}^{+\infty}\frac{\sinh(2kx)}{x\cosh^{2n+2}x}\,dx\quad(k\in\llbracket 1,n\rrbracket)$ $\int_{0}^{\tfrac{\pi}{4}}\frac{\sin(4nx)}{\ln(\tan x)}\,dx$ $\int_{0}^{1}\frac{\operatorname{Li}_{-2n-1}(-x^{2})}{x}\,\ln\ln\tfrac{1}{x}\,dx$ $\int_{0}^{1}\frac{x^{2n-1}}{\operatorname{arctanh}x}\,dx$ $\int_{0}^{+\infty}\frac{\tanh^{2n}x}{x^{2}}\,dx$ $\int_{0}^{+\infty}\frac{\tanh^{\,n+1}x}{x^{\,n+1}}\,dx$ $\tiny{\displaystyle\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}x}\,dx\quad(k\in\llbracket 0,n-1\rrbracket)}$ $\int_{0}^{+\infty}\frac{\sinh(2kx)}{x\cosh^{2n+1}x}\,dx\quad(k\in\llbracket 1,n\rrbracket)$ $\int_{0}^{\tfrac{\pi}{4}}\frac{\cos\!\big((4n-2)x\big)}{\ln(\tan x)}\,dx$ $\int_{0}^{1}\frac{\operatorname{Im}\!\left[\operatorname{Li}_{-2n}(ix)\right]}{x}\,\ln\ln\tfrac{1}{x}\,dx$ $\int_{0}^{1}\frac{x^{2n-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}x}\,dx$
We remark that the integral $\int_{0}^{1}\frac{\operatorname{Im}\!\left[\operatorname{Li}_{-2n}(ix)\right]}{x}\,\ln\ln\tfrac{1}{x}\,dx$ generalizes the integral
$\dfrac{\beta(2)}{\pi}=\displaystyle\int_{0}^{1}\dfrac{u^{4}-6u^{2}+1}{2(1+u^{2})^{3}}\,\ln\ln\frac{1}{u}\,du$
that has been firstly evaluated by Adamchik [5] and rediscovered by Blagouchine [4]. Additionally, the integral $\int_{0}^{1}\frac{\operatorname{Li}_{-2n-1}(-x^{2})}{x}\,\ln\ln\tfrac{1}{x}\,dx$ exhibits a general form of $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ by also extending the Blagouchine’s integral $\int_{0}^{1}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx=\frac{7\,\zeta(3)}{8\,\pi^{2}}$ to all higher $n$. The general formulae and closed-form expressions are provided by the following propositions.
Proposition 2.6
$\forall\,n\in\mathbb{N}^{*},\quad\int_{0}^{1}\frac{\mathrm{Li}_{-2n-1}(-x^{2})\,\ln\ln\frac{1}{x}}{x}\,dx=(-1)^{n}\left(1-\frac{1}{2^{2n+1}}\right)\,\dfrac{(2n)!}{2}\frac{\zeta(2n+1)}{\pi^{2n}}$
Proof
Let $J_{n}:=\displaystyle\int_{0}^{1}\dfrac{\mathrm{Li}_{-2n-1}(-x^{2})\ln\ln\frac{1}{x}}{x}\,dx$. On recalling the definition of the polylogarithm $\mathrm{Li}_{s}(z):=\sum_{k=1}^{\infty}\dfrac{z^{k}}{k^{s}}$ and bearing in mind that $|-x^{2}|<1$ on the domain of integration, one becomes
$J_{n}=\int_{0}^{1}\dfrac{1}{x}\,\displaystyle\sum_{k=1}^{\infty}\dfrac{(-1)^{k}\,x^{2k}}{k^{-2n-1}}\ln\ln\frac{1}{x}\,dx=\int_{0}^{1}\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\,x^{2k-1}k^{2n+1}\ln\ln\frac{1}{x}\,dx$
After switching the order of the operators (assumed as justified), one has
$J_{n}=\displaystyle\sum_{k=1}^{\infty}(-1)^{k}k^{2n+1}\int_{0}^{1}x^{2k-1}\ln\ln\frac{1}{x}\,dx$
And applying the well-known formula $\int_{0}^{1}x^{s}\ln\left(\ln\frac{1}{x}\right)\,dx=-\dfrac{\gamma+\ln(s+1)}{s+1}$ gives
$J_{n}=-\displaystyle\sum_{k=1}^{\infty}(-1)^{k}k^{2n+1}\dfrac{\gamma+\ln(2k)}{2k}=-\dfrac{1}{2}\displaystyle\sum_{k=1}^{\infty}(-1)^{k}k^{2n+1}\dfrac{\ln\left(2e^{\gamma}\right)+\ln k}{k}$
The later sum is then split in two to provide
$J_{n}=-\dfrac{\ln\left(2e^{\gamma}\right)}{2}\underbrace{\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\dfrac{1}{k^{-2n}}}_{\eta(-2n)}-\dfrac{1}{2}\underbrace{\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\dfrac{\ln k}{k^{-2n}}}_{\eta^{\prime}(-2n)}$
One recognizes easily the Dirichlet $\eta$ function and its derivative. Now, the $\eta$-function has zeros at negative even integers since the Riemann $\zeta$ function also has zeros at negative even integers [6][7][8]. And its derivative is
$\eta^{\prime}(s)=\left(1-2^{1-s}\right)\zeta^{\prime}(s)+2^{1-s}\ln 2\cdot\zeta(s)$
. It remains
$J_{n}=-\dfrac{1}{2}\left(1-2^{1+2n}\right)\zeta^{\prime}(-2n)$
And the closed form of the derivative zeta at even negative integers [9][10]
$\zeta^{\prime}(-2n)=(-1)^{n}\,\frac{(2n)!}{2\,(2\pi)^{2n}}\,\zeta(2n+1)$
yields
$J_{n}=(-1)^{n}\left(1-\frac{1}{2^{2n+1}}\right)\,\dfrac{(2n)!}{2}\frac{\zeta(2n+1)}{\pi^{2n}}$
Proposition 2.7
$\forall n\in\mathbb{N}^{*},\int_{0}^{1}\dfrac{\operatorname{Im}\!\bigl[\operatorname{Li}_{-2n}(i\,x)\bigr]}{x}\,\ln\ln\frac{1}{x}\,dx=(-1)^{n+1}\dfrac{2^{2n-1}(2n-1)!}{\pi^{2n-1}}\beta(2n)$
Proof
Malmsten proved in 1842 [4] the following functional equation
$\beta(1-s)=\left(\frac{\pi}{2}\right)^{-s}\sin\left(\frac{\pi s}{2}\right)\Gamma(s)\,\beta(s)$
Let $\chi(s):=\left(\dfrac{\pi}{2}\right)^{-s}\sin\left(\dfrac{\pi s}{2}\right)\Gamma(s)$, so we get $\beta(1-s)=\chi(s)\,\beta(s)$. On differentiating with respect to $s$, we get
$\beta^{\prime}(1-s)=\chi^{\prime}(s)\,\beta(s)+\chi(s)\,\beta^{\prime}(s)$
We see trivially that $\chi(2n)=0$ due to the $\sin$ function. It remains
$\beta^{\prime}(1-2n)=\chi^{\prime}(2n)\,\beta(2n)$
One has
$\chi^{\prime}(2n)=\left.\dfrac{d}{ds}\left(\chi(s)\right)\right|_{s=2n}$
. Let $\Delta(s):=\left(\dfrac{\pi}{2}\right)^{-s}\Gamma(s)$, so that $\chi(s)=\Delta(s)\sin\left(\dfrac{\pi s}{2}\right)$. On differentiating with respect to $s$, we have
$\chi^{\prime}(s)=\Delta^{\prime}(s)\sin\left(\dfrac{\pi s}{2}\right)+\dfrac{\pi}{2}\Delta(s)\cos\left(\dfrac{\pi s}{2}\right)$
On evaluating this equality for $s=2n$ and putting it back in the earlier relation, we obtain
$\beta^{\prime}(1-2n)=(-1)^{n+1}\dfrac{2^{2n-1}(2n-1)!}{\pi^{2n-1}}\beta(2n)\quad(2.8)$
Let $R_{n}:=\displaystyle\int_{0}^{1}\dfrac{\operatorname{Im}\!\bigl[\operatorname{Li}_{-2n}(i\,x)\bigr]}{x}\,\ln\ln\frac{1}{x}\,dx$. One has $\dfrac{\operatorname{Im}\!\bigl[\operatorname{Li}_{-2n}(i\,x)\bigr]}{x}=\displaystyle\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n}x^{2k}$
The Malmsten’s integral becomes
$R_{n}=\int_{0}^{1}\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n}x^{2k}\,\ln\ln\frac{1}{x}\,dx=\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n}\int_{0}^{1}x^{2k}\,\ln\ln\frac{1}{x}\,dx$
The interchange of summation and integration being assumed as correct. On evaluating the integral, we get
$\int_{0}^{1}x^{2k}\,\ln\ln\frac{1}{x}\,dx=-\dfrac{\gamma+\ln(2k+1)}{2k+1}$
We become
$R_{n}=-\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n}\dfrac{\gamma+\ln(2k+1)}{2k+1}$
Which reads
$R_{n}=-\gamma\underbrace{\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n-1}}_{\beta(1-2n)}-\underbrace{\sum_{k=0}^{\infty}(-1)^{k}\,(2k+1)^{2n}\dfrac{\ln(2k+1)}{2k+1}}_{-\beta^{\prime}(1-2n)}$
One reads easily from the functional equation of $\beta$ that $\beta(1-2n)=0$. The value of $\beta^{\prime}(1-2n)$ follows from (2.8).
We conclude
$R_{n}=\beta^{\prime}(1-2n)=(-1)^{n+1}\dfrac{2^{2n-1}(2n-1)!}{\pi^{2n-1}}\beta(2n)$
§3. Sketch of Methods
All integrals recorded in the above table without exceptions can be reduced to linear combinations of
$\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{n}\!x}\,dx$
with rational coefficients. The following lines only give instructions on how to prove that
$\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$
holds form (2.4) and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ holds form (2.5)
The study of the integrals $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ relays solely on contour integration techniques. We made the remark that the Laurent expansion approach is very fruitful to find the residues here. To find these residues, we note that $\cosh^{n}\!z$ has a pole of $n$-th order at $z_{l}:=\dfrac{(2l+1)i\pi}{2}$ where $l\in\mathbb{Z}$. We find the Laurent expansion of $\dfrac{1}{\cosh^{n}\!z}{}$ around $z_{l}$ by letting first of all $z=z_{l}+w$. It comes
$\cosh^{n}\!z=(\cosh z_{l}\,\cosh w+\sinh z_{l}\,\sinh w)^{n}=\sinh^{n}z_{l}\,\sinh^{n}w$
since $\cosh z_{l}=0$. Hence, finding the Taylor expansion of $\cosh^{n}\!z$ around $z_{l}$ reduces to finding the Taylor expansion of $\sinh^{n}z_{l}\,\sinh^{n}w$ around $w=0$, which is done after linearizing $\sinh^{n}w$ depending of the parity of $n$ and using the Taylor expansion of $\sinh(\alpha w)$ around $w=0$. With the Cauchy’s product formula we exhibit the Laurent expansion of $\cosh^{-n}\!z$ around $z_{l}$ and remark rational coefficients appearing in these series. On noticing additionally that $\dfrac{\sinh\!\left((2k+1)z\right)}{z}$ is analytic around $z=z_{l}$, we find its Taylor expansion by using the Leibniz’s formula for derivatives of products on bearing in mind these formulae which are quite easy to validate through induction on $n$:
$\dfrac{d^{n}}{dx^{n}}(x^{-1})=(-1)^{n}\,n!\,x^{-(n+1)}$
$\frac{d^{n}}{dx^{n}}\sinh(ax)=a^{n}\times\begin{cases}\cosh(ax)&\text{if }n\text{ is odd},\\ \sinh(ax)&\text{if }n\text{ is even}.\end{cases}$
The residue is read after multiplying the Laurent expansion of $\cosh^{-n}\!z$ by the Taylor expansion of $\dfrac{\sinh\!\left((2k+1)z\right)}{z}$. The contour of integration, designated by $C_{N,r}$ is the rectangle bounding the region of the complex plane defined by $\begin{cases}0\leq\,Im(z)\leq 2N\pi\,\,(N\in\mathbb{N})\\ \left|Re\,(z)\right|\leq r\,\,(r\in\mathbb{R_{+}^{*}})\end{cases}$. The modulus of the integrand vanishes on the paths $\begin{cases}0\leq\,Im(z)\leq 2N\pi\,\,\\ \left|Re\,(z)\right|=r\,\,\end{cases}$ and on the axis $\begin{cases}Im(z)=2N\pi\,\,\\ \left|Re\,(z)\right|\leq r\,\,\end{cases}$ for sufficiently large $N$ and $r$. Cauchy’s residue theorem confirms the claim at the end.
The proofs of Malmsten’s integrals outlined in Propositions 2.6 and 2.7 lack full mathematical rigor for two main reasons. First, the interchange of summation and integration is not properly justified. Second, the manipulations involve regularized versions of the beta and zeta functions, whose corresponding real-valued expressions would otherwise diverge. To establish a rigorous proof of these formulae, it is therefore necessary to derive explicit expressions for the coefficients $L_{p,n}$ and $K_{p,n}$ satisfying the given relations :
$\int_{0}^{1}\frac{\mathrm{Li}_{-2n-1}(-x^{2})\,\ln\ln\frac{1}{x}}{x}\,dx=\sum_{p=1}^{n}L_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}\quad\text{and }\quad\int_{0}^{1}\dfrac{\operatorname{Im}\!\bigl[\operatorname{Li}_{-2n}(i\,x)\bigr]}{x}\,\ln\ln\frac{1}{x}\,dx=\sum_{p=1}^{n}K_{p,n}\,\frac{\beta(2p)}{\pi^{2p-1}}$
, to compute the coefficients $L_{n,n}$ and $K_{n,n}$ and to show that
$L_{p,n}=K_{p,n}=0\qquad\forall p\in\llbracket 1,n-1\rrbracket$
This method is fully explored and detailled in [3]. It starts by showing first of all the following identities :
[…]
§4. Conjectures and Perspectives
In this paragraph, we mean by $\widetilde{\beta\zeta}(n)$ either the number $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ or $\dfrac{\beta(2n)}{\pi^{2n-1}}$ with $n\in\mathbb{N}^{*}$
Proposition 4.1
Let $\mathbb{K}$ be a field, $V$ a $\mathbb{K}$-vector space, $I_{n},J_{n}$ two sequences in $V$ such that holds :
$\forall n\in\mathbb{N}^{*},\exists(x_{1,n},x_{2,n},\ldots,x_{n,n})\in\mathbb{K}^{n}:I_{n}=\sum_{k=1}^{n}x_{k,n}J_{k}.$
Then
$\forall n\in\mathbb{N}^{*},\exists(y_{1,n},y_{2,n},\ldots,y_{n,n})\in\mathbb{K}^{n}:J_{n}=\sum_{k=1}^{n}y_{k,n}I_{k}$
Proof :
One has:
$I_{1}=x_{1,1}J_{1}$
$I_{2}=x_{1,2}J_{1}+x_{2,2}J_{2}$
$\vdots$
$I_{n}=x_{1,n}J_{1}+x_{2,n}J_{2}+\ldots+x_{n,n}J_{n}$
The last exercise consists of expressing each $J_{k}$ as linear combination of $I_{k}$ which corresponds to the last step of Gauss elimination algorithm.
Corollary : Since there exist integrals $I_{n}$ either holding form (2.4) or (2.5), we can always find rational coefficients $y_{k,n}$ satisfying $\widetilde{\beta\zeta}(n)=\displaystyle\sum_{k=1}^{n}y_{k,n}I_{k}$. The irrationality of $\dfrac{\widetilde{\beta\zeta}(n)}{\pi}$ would be deduced then from this conjecture :
$\pi\notin\text{Span}_{\mathbb{Q}}\left\{I_{1},\ I_{2},\ I_{3},\ I_{4},\ I_{5},\ I_{6},\ I_{7},\ I_{8},\ I_{9},\ I_{10},\ I_{11},\ I_{12},\ \ldots\right\}$
Application to conjectures :
  • From the table of paragraph §2 and in combination with proposition 4.1, we conclude the existence of rational coefficients $\chi_{p,n}$ and $\xi_{p,n}$ which satisfy
    $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}\chi_{p,n}\displaystyle\int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{2p+1}x}\,dx\quad\text{and }\quad\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}\xi_{p,n}\displaystyle\int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{2p}x}\,dx$
    The irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ would be deduced from this lonely following conjecture:
    $\pi\notin\text{Span}_{\mathbb{Q}}\left\{\int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{2}x}\,dx,\ \int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{3}x}\,dx,\ \int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{4}x}\,dx,\ \int_{0}^{+\infty}\frac{\sinh x}{x\,\cosh^{5}x}\,dx,\ \ldots\right\}$
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}_{,}x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $r_{p,n}$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}r_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}_{,}x}\,dx$; thus, the existence of polynomials $\Xi_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\int_{0}^{1}\frac{x\,\Xi_{n}(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx$.
    Few first polynomials $\Xi_{n}$ are [3]:
    $\Xi_{1}(x)=\frac{1}{4}$
    $\Xi_{2}(x)=-\frac{1}{16}\,x^{2}+\frac{5}{96}$
    $\Xi_{3}(x)=\frac{1}{64}\,x^{4}-\frac{3}{128}\,x^{2}+\frac{61}{7\,680}$
    $\Xi_{4}(x)=-\frac{1}{256}\,x^{6}+\frac{13}{1536}\,x^{4}-\frac{173}{30\,720}\,x^{2}+\frac{277}{258\,048}$
[…]
    On multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\beta(2n)}{\pi^{2n}}$ for each $n$ would follow from the following conjecture:
    $\forall P\in\mathbb{Z}[x]\backslash\left\{0\right\},\displaystyle\int_{0}^{1}\frac{x\,P(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx\notin\pi\mathbb{Q}$
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}_{,}x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $s_{p,n}$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}s_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}_{,}x}\,dx$; thus, the existence of polynomials $\Lambda_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\int_{0}^{1}\frac{x\,\Lambda_{n}(x)}{\text{arctanh }x}\,dx$.
    Few first polynomials $\Lambda_{n}$ are [3]:
    $\Lambda_{1}(x)=\frac{1}{7}$
    $\Lambda_{2}(x)=-\frac{1}{31}\,x^{2}+\frac{2}{93}$
    $\Lambda_{3}(x)=\frac{1}{127}\,x^{4}-\frac{4}{381}\,x^{2}+\frac{17}{5\,715}$
    $\Lambda_{4}(x)=-\frac{1}{511}\,x^{6}+\frac{2}{511}\,x^{4}-\frac{6}{2\,555}\,x^{2}+\frac{62}{160\,965}$
[…]
    On multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ for each $n$ would follow from the following conjecture:
    $\forall P\in\mathbb{Z}[x]\backslash\left\{0\right\},\displaystyle\int_{0}^{1}\frac{x\,P(x)}{\text{arctanh }x}\,dx\notin\pi\mathbb{Q}$
In the interest of the theory of Fourier series :
Since proposition 4.1 allows to conclude that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\sum_{p=1}^{n}r_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4px)}{\ln(\tan x)}\,dx$ and that
$\forall n\in\mathbb{N}^{*},\dfrac{\beta(2n)}{\pi^{2n-1}}=\sum_{p=1}^{n}\Psi_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\cos((4p-2)x)}{\ln(\tan x)}\,dx$
, with $r_{p,n},\Psi_{p,n}\in\mathbb{Q}$, the irrationality of the numbers $\dfrac{\beta(2n)}{\pi^{2n}}$ and $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ would be deduced from this conjecture :
$\pi\notin\text{Span}_{\mathbb{Q}}\left\{\int_{0}^{\frac{\pi}{4}}\frac{\cos(2x)}{\ln(\tan x)}\,dx,\ \int_{0}^{\frac{\pi}{4}}\frac{\sin(4x)}{\ln(\tan x)}\,dx,\ \int_{0}^{\frac{\pi}{4}}\frac{\cos(6x)}{\ln(\tan x)}\,dx,\ \int_{0}^{\frac{\pi}{4}}\frac{\sin(8x)}{\ln(\tan x)}\,dx,\ \ldots\right\}$
Furthermore, we introduce a new vector space. The integrals $\int_{0}^{\frac{\pi}{4}}\frac{\sin(4px)}{\ln(\tan x)}\,dx$ and $\int_{0}^{\frac{\pi}{4}}\frac{\cos((4p-2)x)}{\ln(\tan x)}\,dx$ evaluated in the previous sections convince us to introduce the study of the integral $\int_{0}^{\frac{\pi}{4}}\dfrac{f(x)}{\ln(\tan x)}\,dx$ where $f$ is a function of the real variable. This integral converges if and only if
$\exists\,m\geq 1:\lim_{x\rightarrow\frac{\pi}{4}}\dfrac{f(x)}{(x-\frac{\pi}{4})^{m}}\in\mathbb{R}$
In order words, the integral $\int_{0}^{\frac{\pi}{4}}\dfrac{f(x)}{\ln(\tan x)}\,dx$ converges if and only if $f$ has a root at $x=\dfrac{\pi}{4}$ with multiplicity at least 1.
We consider in the same spirit this trivial theorem
$\int_{0}^{\frac{\pi}{4}}\dfrac{f(x)}{\ln(\tan x)}\,dx$ is a rational multiple of $\pi$ if and only if there exists a function $g$ so that holds :
[…]
We remind now that
$\frac{\zeta(2n+1)}{\pi^{2n}}=\sum_{p=1}^{n}r_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4px)}{\ln(\tan x)}\,dx\quad\text{and}\quad\displaystyle\frac{\beta(2n)}{\pi^{2n-1}}=\sum_{p=1}^{n}s_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\cos((4p-2)x)}{\ln(\tan x)}\,dx$
, with $r_{p,n},s_{p,n}\in\mathbb{Q}^{*}$. Swapping the integrator and the sum yields
[…]
Instinctively, the functions $h_{n}(x):=\displaystyle\sum_{p=1}^{n}r_{p,n}\sin(4px)$ and $k_{n}(x):=\displaystyle\sum_{p=1}^{n}s_{p,n}\cos((4p-2)x)$ cannot be written in the form $g(x)\ln(\tan x)$ with a normal common function $g$ until
$g(x):=\dfrac{\psi(x)\displaystyle\sum_{p=1}^{n}r_{p,n}\sin(4px)}{\ln(\tan x)}\quad\text{or}\quad g(x):=\dfrac{\psi(x)\displaystyle\sum_{p=1}^{n}s_{p,n}\cos((4p-2)x)}{\ln(\tan x)}$
where $\psi$ is a normal common function. The rationality of the ratios $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ would be surprising. And the same line of reasoning holds for $\dfrac{\beta(2n)}{\pi^{2n}}$
To address the irrationality (respectively, transcendence) of the ratios $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ in a rigorous and conclusive way, by pursuing only one idea, we propose to study the $\mathbb{Q}$-vector space of functions $f$ such that
$\int_{0}^{\frac{\pi}{4}}\frac{f(x)}{\ln(\tan x)}\,dx\in\mathbb{Q}\pi\quad\left(\text{respectively}\quad\int_{0}^{\frac{\pi}{4}}\frac{f(x)}{\ln(\tan x)}\,dx\in\overline{\mathbb{Q}}\pi\right),$
and to demonstrate that these functions possess at least one structural property neither shared by finite trigonometric sums of the form
$h_{n}(x):=\sum_{p=1}^{n}a_{p,n}\sin(4px),\quad\text{with }a_{p,n}\in\mathbb{Z}.$
nor
$k_{n}(x):=\sum_{p=1}^{n}b_{p,n}\cos((4p-2)x),\quad\text{with }b_{p,n}\in\mathbb{Z}.$
The last theorem ensures the existence of the functions being the solution of the equation but only their properties interest us here. For instance, they may admit an infinite Fourier expansion. We consider the form of $f$ under two cases :
  • Either $f$ has the form $\sum_{p=1}^{\infty}\left[a_{p}\sin(4px)+b_{p}\cos((4p-2)x)\right]$ where at least one of the sets $\left\{a_{p}\right\}$ and $\left\{b_{p}\right\}$ holds infinite non zero values;
  • Or $f$ has the form $\sum_{p=1}^{N}\left[a_{p,N}\sin(4px)+b_{p,N}\cos((4p-2)x)\right]$ where both sets $\left\{a_{p,N}\right\}$ and $\left\{b_{p,N}\right\}$ contains at least a non zero coefficient, with $N$ natural integer.
This reformulation of the problem is independent of any particular value of $n$, yet it is robust enough to imply the irrationality (respectively, transcendence) of both the ratios $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ for all integers $n\geq 1$.
In fact, the focus of the analysis should be the integral
$\int_{0}^{\frac{\pi}{4}}\frac{f(x)}{\ln(\tan x)}\,dx,$
especially to exhibit when it yields a rational (respectively, algebraic) multiple of $\pi$. It is very likely that there does not exist a function $g\in L^{1}\left([0,\tfrac{\pi}{4}]\right)$ such that
$h_{n}(x)\quad or\quad k_{n}(x)=g(x)\ln(\tan x)\quad\text{and}\quad\int_{0}^{\frac{\pi}{4}}g(x)\,dx\in\mathbb{Q}\pi.$
However, a rigorous proof of this claim remains to be established.
We establish an interesting identity, which lends positive support to the validity of this conjecture. Of course we recognize the weakness and lack of rigor of our lines of reasoning, because of the convergence and summation/integration swap. We start from the well-known Fourier series expansion of $\ln\tan x$
$\ln(\tan x)\;=\;-2\sum_{k=0}^{\infty}\frac{\cos\!\bigl((4k+2)x\bigr)}{2k+1},\qquad 0<x<\tfrac{\pi}{2}.$
On dividing each side by $\ln\tan x$ -despite of not rigorous justification of convergence, one has
$1\;=\;-2\sum_{k=0}^{\infty}\frac{\cos\!\bigl((4k+2)x\bigr)}{(2k+1)\,\ln(\tan x)},\qquad 0<x<\tfrac{\pi}{2}.$
On integrating terms by terms -despite lack of rigor, we obtain the formal identity
$\frac{\pi}{4}\;=\;-2\sum_{k=0}^{\infty}\frac{1}{2k+1}\int_{0}^{\pi/4}\frac{\cos\!\bigl((4k+2)x\bigr)}{\ln(\tan x)}\,dx.$
Some computations of the first few partial sums on the right handside series tend to show that this identity may be correct. This identity does not stay in coherence with the linear dependance of the previous set stated in our last conjecture.